Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Answer by rapaljer(4667)

Let x= number of nickels @ 5 cents
50-x = number of dimes @ 10 cents

The equation is
5x+10(50-x) = 350 cents
5x+ 500-10x=350
-5x=350-500
-5x=-150
x=30 Nickels
50-30= 20 Dimes

Check:
.05(30)+.10(20)=
1.50+2.00
$3.50

For additional explanations on Word Problems, especially coin problems, please see my website by clicking on my tutor name "rapaljer" anywhere in algebra.com. Click on the first link, which is "Basic and Intermediate Algebra". Look for Basic Algebra, Chapter 1, Section 1.09 and 1.10. For solutions in "LIVING COLOR", look for my "MATH IN LIVING COLOR" page, and click on Basic Algebra, Chapter 1, Sections 1.10. The explanations are written for people who need help with math!

R^2